Calculate the molarity of a solution containing 5g of NaOH in 450 mL solution

Grade:12

2 Answers

Sudhanshu Shekhar
5 years ago
molarity= moles of solute/ vol.of soln
 
Moles = given mass / mol.mass
            = 5/40
            =1/8
Now , 
      Molarity= 1/8×450×10*-3
                    =0.27m
Mayank Raj
5 years ago
Molarity=moles of solute/vol of solution in litre
Moles of solute =5/40=1/8 (Mw of NAOH=40)
MOLARITY =1/8*1000/450
=0.27M